Financial Statements of a Manufacturing Firm
The following events took place for Rushmore Biking Inc. during February, the first month of operations as a producer of road bikes:

Purchased $480,000 of materials.

Used $434,500 of direct materials in production.

Incurred $125,000 of direct labor wages.

Applied factory overhead at a rate of 40% of direct labor cost.

Transferred $578,000 of work in process to finished goods.

Sold goods with a cost of $550,000.

Revenues earned by selling bikes, $910,000.

Incurred $185,000 of selling expenses.

Incurred $90,000 of administrative expenses.
a. Prepare the income statement for Rushmore Biking Inc. for the month ending February 28. Assume that Rushmore Biking Inc. uses the perpetual inventory method.
Rushmore Biking Inc.
Income Statement
For the Month Ended February 28

Revenues

$

Cost of goods sold

Gross profit

$
Selling and administrative expenses:

Selling expenses


$


Administrative expenses


Total selling and administrative expenses


Operating income

$


b. Determine the inventory balances on February 28, the end of the first month of operations.
Materials inventory, February 28
$
Work in process inventory, February 28
$
Finished goods inventory, February 28
$
Answer #1

a. Prepare the income statement for Rushmore Biking Inc. for the month ending February 28.

Assume that Rushmore Biking Inc. uses the perpetual inventory method.

Rushmore Biking Inc
Income Statement
For the Month Ended February 28
Revenues $910000
Cost of goods sold -550000
Gross profit $360000
Selling and administrative expenses:
Selling expenses $185000
Administrative expenses 90000
Total selling and administrative expenses -275000
Operating income $85000

b. Determine the inventory balances on February 28, the end of the first month of operations.

Materials inventory, February 28= Purchases of materials-Used of direct materials in production

= $480000-434500= $45500

Work in process inventory, February 28= Used of direct materials in production+Direct labor wages incurred+Factory overhead applied-Transferred to finished goods

= $434500+125000+($125000*40%)-578000= $31500

Finished goods inventory, February 28= Transferred to finished goods-Cost of goods sold

= $578000-550000= $28000